# Batched Nonparametric Contextual Bandits

### v2

@@ 1.1 Main contributions
  • In addition, we demonstrate that the aforementioned lower bound is tight by providing a matching upper bound (up to log factors). Specifically, we design a novel algorithm—Batched Successive Elimination with Dynamic Binning (BaSEDB)—for the nonparametric bandits with batch constraints. BaSEDB progressively splits the covariate space into smaller bins whose widths are carefully selected to align well with the corresponding batch size. The delicate interplay between the batch size and the bin width is crucial for obtaining the optimal regret in the batch setting.
@@ 1.2 Related work
The batch constraint has received increasing attention in recent years. [40, 21] considered the multi-armed bandit problem under the batch setting and showed that $O(\log\log T)$ batches are adequate in achieving the rate-optimal regret, compared to the fully online setting. [26, 47] extended batch learning to the (generalized) linear contextual bandits and [46, 56, 17] further studied the setting with high-dimensional covariates. [29, 28] established batch learning guarantees for the Thompson sampling algorithm. [18] considered Lipschitz continuum-armed bandit problem with the batch constraint. Inference for batched bandits was considered in [60]. A concept related to batch learning in literature is called delayed feedback [14, 13, 55, 19]. These works consider the setting where rewards are observed with delay and analyze effects of delay on the regret. [32, 2] studied delayed feedback in nonparametric bandits and the key difference to batch learning is that the batch size is given, whereas in our case, it is a design choice by the statistician. Batch learning’s focus is different to that of delayed feedback in the sense that the former gives the decision maker discretion to choose the batch size which makes it possible to approximate the optimal standard online regret with a small number of batches. Finally, the notion switching cost is intimately related to the batch constraint. [12] studied online learning with low switching cost and obtained minimax optimal regret with $O(\log\log T)$ batches. [5, 61, 20, 57, 44] developed regret guarantees with low switching cost for reinforcement learning. Low switching cost can be interpreted as infrequent policy updates, but it does not require the learner to divide the samples into batches with feedback only becoming available at the end of a batch.
@@ 2 Problem setup
The goal of the statistician is to design an $M$-batch policy $(\Gamma,\pi)$ that can compete with an oracle that has perfect knowledge (i.e., the law of $(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)})$) of the environment. Formally, we define the cumulative regret as
@@ 2.1 Assumptions
Throughout the paper, we assume that $\alpha\beta\leq 1$. By proposition 2.1 from [48], when $\alpha\beta>1$, one of the arms will dominate the other one for the entire covariate space. The instance is reduced to a multi-armed bandit without covariates which is not the interest of the current paper. Therefore, we focus on the case $\alpha\beta\leq 1$ hereafter.
@@ 3 Fundamental limits of batched nonparametric bandits
Recall that $\mathcal{F}(\alpha,\beta)$ denotes the class of nonparametric bandit instances (i.e., distributions over (1)) that obey Assumptions 1-2. We have the following minimax lower bound for any $M$-batch policy with a fixed grid, in which we define
@@
$\mathbb{E}[R_{T}(\pi)]\geq\tilde{D}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
See Section 3.1 for the proof of this lower bound.
As a sanity check, one sees that as $M$ increases, the lower bound decreases. This is intuitive, as the policy is more powerful as $M$ increases. As a result, the problem of batched nonparametric bandits becomes easier.
@@ 3.2 Lower bound for general $M$-batch policy
$\mathbb{E}[R_{T}(\pi)]\geq\tilde{D}_{1}(\frac{1}{M})^{\tilde{D}_{2}}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@ 4.1 Regret guarantees
$\mathbb{E}[R_{T}(\hat{\pi})]\leq\tilde{C}(\log T)^{2}\cdot T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
$\mathbb{E}[R_{T}(\hat{\pi})]\leq\tilde{C}(\log T)^{2}\cdot T^{1-\gamma},$
@@ 4.3 Failure of static binning
$\mathbb{E}[R_{T}(\hat{\pi}_{\mathrm{static}})]\geq\tilde{C}_{1}T^{\frac{9}{19}+\kappa},$
@@ 5 Adaptivity to margin parameter $\alpha$
In this section, we provide some discussions on the possibility of adapting to the margin parameter $\alpha$ if it is unknown. Recall in Section 4, the grid choice of Algorithm 1 requires knowledge of $\alpha$. One may ask if such knowledge is essential in obtaining small regret. Unfortunately, the following theorem demonstrates that the price of not knowing $\alpha$ is at least a polynomial increase in regret.
@@
Consider $M=2$ (or $3$) and $\beta=d=1$. For any algorithm that does not know the true margin parameter $\alpha^{\star}$, there exists a choice of $\alpha^{\star}$ such that
$\sup_{\mathcal{F}(\alpha^{\star},1)}\mathbb{E}[R_{T}(\pi)]\geq\tilde{D}_{3}T^{\frac{1-\frac{\alpha^{\star}+1}{3}}{1-(\frac{\alpha^{\star}+1}{3})^{M}}+\kappa_{1}},$
for some $\tilde{D}_{3}>0,\kappa_{1}$ that are independent of $T$.
@@
Theorem 5 says for any algorithm that does not have knowledge of $\alpha^{\star}$, its regret is at least a polynomial factor larger than the optimal regret attained by Algorithm 1. This result shows batch learning for nonparametric bandits is much harder than the fully online case to some extent, where adaptivity to $\alpha^{\star}$ could be achieved for free [39].
The intuition behind the proof of Theorem 5 is that since the algorithm does not know $\alpha^{\star}$, it has little hope to pick the first batch size $t_{1}$ optimally. If $t_{1}$ is too large, then the adversary can choose a big $\alpha^{\star}$, which corresponds to the family of reward functions with larger gaps, so that the algorithm’s regret during the first batch explodes. If $t_{1}$ is too small, then the adversary can choose a little $\alpha^{\star}$, which corresponds to the family of reward functions with smaller gaps, so that the algorithm’s knowledge gathered during the first batch is not enough to distinguish the arms and its regret will explode in later batches.
6 Conclusions
@@
Adapting to margin parameters.
While we have shown that adaptivity to the margin parameter is not possible when the batch constraint is stringent, i.e., when $M=2$ (or 3), it leaves open the question of designing optimal adaptive algorithm when $M$ is large, say $M\asymp\log\log T$. In fact, when $M=T$, i.e., in the fully online setting, [39] provides an adaptively binned successive elimination algorithm that is capable of adapting to the margin parameter optimally.
@@ Appendix A Proof of Theorem 2
Define $b\asymp T^{(1-\gamma)/(1-\gamma^{M})}$. For each $1\leq i\leq M$, we set $T_{i}=\lfloor b^{(1-\gamma^{i})/(1-\gamma)}\rfloor$, $z_{i}=\lceil(36T_{i-1}M^{2})^{1/(2\beta+d)}\rceil$, and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. We reuse the family of hard instances $\mathcal{C}_{z_{i}}$ as defined in (4), and define the mixture distribution
@@
where $q_{i}:\Omega_{s_{i}}\rightarrow[0,1]$ is selected so that the above equality holds, $\mathbb{P}_{\pi,i,f_{\omega}}$ is the distribution of the observations when $f_{\omega}\in\mathcal{C}_{z_{i}}$. It is easy to see $\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)=1$.
We pause here to state a useful claim regarding the family $\{Q_{i}\}_{i=1}^{M}$ of mixture distributions, namely, they are all close to each other under the total variation distance.
@@
Define the event
$A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}.$
Intuitively, $A_{i}$ models the event when the algorithm’s selected grid points $t_{i-1}$ and $t_{i}$ are suboptimal. When $A_{i}$ happens, the goal is to design a problem instance such that using observations up $t_{i-1}$ cannot distinguish the optimal arm and therefore the policy must incur a large regret between $t_{i-1}$ and $t_{i}$.
The following lemma ensures that at least one of the bad events $A_{i}$’s happens with sufficiently large probability under the mixture distribution.
@@
There exists $i^{\star}\in[M]$ such that $Q_{i^{\star}}(A_{i^{\star}})\geq 1/(2M)$.
@@
If $Q_{i}(A_{i})\geq 1/(2M)$, then
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}z_{i}^{-\beta(1+\alpha)}M^{-\frac{1+\alpha}{\alpha}}.$
Now we are ready to establish the desired claim in the theorem. It is straightforward to see that
@@ A.2 Proof of Lemma 3
For any $1\leq i\leq M$, we have
$|Q_{M}(A_{i})-Q_{i}(A_{i})|\overset{\mathrm{(i)}}{=}|Q_{M}^{T_{i-1}}(A_{i})-Q_{i}^{T_{i-1}}(A_{i})|\overset{\mathrm{(ii)}}{\leq}\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\overset{\mathrm{(iii)}}{\leq}\frac{1}{2M},$ (15)
@@
$\sum_{i=1}^{M}Q_{i}(A_{i})$ $=Q_{M}(A_{M})+\sum_{i=1}^{M-1}Q_{i}(A_{i})$
$=Q_{M}(A_{M})+\sum_{i=1}^{M-1}(Q_{i}(A_{i})-Q_{M}(A_{i})+Q_{M}(A_{i}))$
$\overset{\mathrm{(iv)}}{\geq}Q_{M}(A_{M})+\sum_{i=1}^{M-1}(Q_{M}(A_{i})-\frac{1}{2M})\geq\sum_{i=1}^{M}Q_{M}(A_{i})-\frac{1}{2}\overset{\mathrm{(v)}}{=}\frac{1}{2},$
where step (iv) uses inequality (15), and step (v) uses the fact that $\sum_{i=1}^{M}Q_{M}(A_{i})=1$.
@@ A.3 Proof of Lemma 4
$=\frac{1}{2^{s_{i}}}\cdot\frac{T_{i}}{z_{i}^{d}}\left(s_{i}2^{s_{i}-1}Q_{i}(A_{i})-\frac{3}{2}s_{i}2^{s_{i}-1}\sqrt{\frac{1}{36M^{2}}}\right)$
$\geq\frac{1}{8}T_{i}z_{i}^{-\alpha\beta}\frac{1}{M},$
where the last inequality uses the assumption $Q_{i}(A_{i})\geq 1/(2M)$. Therefore, we arrive at
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}^{-\frac{1}{\alpha}}\left[\sup_{f\in\mathcal{C}_{z_{i}}}S_{t_{i}}(\pi;f)\right]^{\frac{1+\alpha}{\alpha}}\apprge T_{i}z_{i}^{-\beta(1+\alpha)}M^{-\frac{1+\alpha}{\alpha}}.$

### v3

@@ 1.1 Main contributions
  • In addition, we demonstrate that the aforementioned lower bound is tight by providing a matching upper bound (up to log factors). Specifically, we design a novel algorithm—Batched Successive Elimination with Dynamic Binning (BaSEDB)—for the nonparametric bandits with batch constraints. BaSEDB progressively splits the covariate space into smaller bins whose widths are carefully selected to align well with the corresponding batch size. The delicate interplay between the batch size and the bin width is crucial for obtaining the optimal regret under the batch setting.
@@ 1.2 Related work
The batch constraint has received increasing attention in recent years. [42, 21] considered the multi-armed bandit problem under the batch setting and showed that $O(\log\log T)$ batches are adequate in achieving the rate-optimal regret, compared to the fully online setting. [26, 49] extended batch learning to the (generalized) linear contextual bandits and [48, 60, 17] further studied the setting with high-dimensional covariates. [31, 30] established batch learning guarantees for the Thompson sampling algorithm. [18] considered Lipschitz continuum-armed bandit problem with the batch constraint. Inference for batched bandits was considered in [64]. A concept related to batch learning in literature is called delayed feedback [14, 13, 59, 19]. These works consider the setting where rewards are observed with delay and analyze effects of delay on the regret. [34, 2] studied delayed feedback in nonparametric bandits and the key difference to batch learning is that the batch size is given, whereas in our case, it is a design choice by the statistician. Batch learning’s focus is different to that of delayed feedback in the sense that the former gives the decision maker discretion to choose the batch size which makes it possible to approximate the optimal standard online regret with a small number of batches. The asymptotic regret of batch learning has been studied in [28, 29] and it is shown that the asymptotic optimality can be achieved with a constant number of batches for multi-armed bandits when the time horizon is known.
Switching cost.
The notion switching cost is intimately related to the batch constraint. [12] studied online learning with low switching cost and obtained the rate-optimal regret with $O(\log\log T)$ switches. [52] established the minimax optimal regret versus policy switch tradeoff in linear bandits with adverserial and stochastic contexts. [5, 65, 20, 61, 46] developed regret guarantees with low switching cost for reinforcement learning. [56] examined collaborative best arm identification under limited interaction, achieving nearly tight tradeoffs between the number of rounds and speedup in multi-armed bandits. Low switching cost can be interpreted as infrequent policy updates, but it does not necessarily require the learner to divide the samples into batches with feedback only becoming available at the end of a batch.
@@ 2 Problem setup
The goal of the statistician is to design an $M$-batch policy $(\Gamma,\pi)$ that can compete with an oracle that has perfect knowledge (i.e., the law of $(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)})$) of the environment. Formally, we define the expected cumulative regret as
@@ 2.1 Assumptions
Throughout the paper, we assume that $\alpha\beta\leq 1$. By proposition 2.1 from [50], when $\alpha\beta>1$, the problem reduces to a static multi-armed bandit where one arm is always optimal, regardless of the context. Since our focus is on contextual bandits, we consider the case $\alpha\beta\leq 1$ hereafter.
@@ 3 Fundamental limits of batched nonparametric bandits
Recall that $\mathcal{F}(\alpha,\beta)$ denotes the class of nonparametric bandit instances (i.e., distributions over (1)) that obey Assumptions 1-2. Compared to multi-armed bandits without covariates, the presence of contexts and the smooth function class no longer allows one to simply decompose the regret as the product between the expected number of times a suboptimal arm is pulled and the suboptimality gap. A more nuanced notion of expected number of suboptimal arm pulls for bandits with covariates was introduced in [50], and its relation to the regret was characterized via an elegant inequality. Nevertheless, a direct application of this relation to the batched setting would lead to suboptimal rates. To achieve a sharp result, we need to construct different families of hard instances for each batch and carefully lower bound the regret within it. We have the following minimax lower bound for any $M$-batch policy with a fixed grid, in which we define
@@
$R_{T}(\pi)\geq\tilde{D}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
Later, we will show that this lower bound is attained (modulo log factors) by our proposed algorithm Batched Successive Elimination with Dynamic Binning (BaSEDB). Hence $T^{\frac{1-\gamma}{1-\gamma^{M}}}$ is the minimax optimal regret for the batched nonparametric bandit problem. See Section 3.1 for the proof of this lower bound.
As a sanity check, one sees that as $M$ increases, the lower bound decreases. This is intuitive, as the policy is more powerful as $M$ increases. As a result, the problem of batched nonparametric bandits becomes easier. When $M=\Omega(\log\log T)$, we recover the minimax regret $\Theta(T^{1-\gamma})$ in the fully online setting.
@@ 3.2 Lower bound for general $M$-batch policy
$R_{T}(\pi)\geq\tilde{D}_{1}(\frac{1}{M})^{\tilde{D}_{2}}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@ 4.1 Regret guarantees
$R_{T}(\hat{\pi})\leq\tilde{C}(\log T)^{2}\cdot T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
$R_{T}(\hat{\pi})\leq\tilde{C}(\log T)^{2}\cdot T^{1-\gamma},$
@@ 4.3 Failure of static binning
$R_{T}(\hat{\pi}_{\mathrm{static}})\geq\tilde{C}_{1}T^{\frac{9}{19}+\kappa},$
@@ 5 Adaptivity to margin parameter $\alpha$
In this section, we provide some discussions on the possibility of adapting to the margin parameter $\alpha$ if it is unknown. Recall in Section 4, the grid choice of Algorithm 1 requires knowledge of $\alpha$. One may ask if such knowledge is essential in obtaining small regret. Unfortunately, the following theorem demonstrates that the price of not knowing $\alpha$ is at least a polynomial increase in regret. Denote $\gamma(\alpha^{\star})=\beta(\alpha^{\star}+1)/(2\beta+d)=\gamma^{\star}$.
@@
Consider $\beta=d=1$. For any algorithm that does not know the true margin parameter $\alpha^{\star}$, there exists a choice of $\alpha^{\star}$ such that
$\sup_{\mathcal{F}(\alpha^{\star},1)}R_{T}(\pi)\geq\tilde{D}_{3}T^{\frac{1-\gamma^{\star}}{1-(\gamma^{\star})^{M}}\;+\;\kappa_{1}}\cdot(\frac{1}{M})^{\tilde{D}_{4}},$
for some $\tilde{D}_{3},\tilde{D}_{4},\kappa_{1}>0$ that are independent of $T$.
@@
Theorem 5 is mostly interesting when the number $M$ of batches is small. In particular, when $M$ is a constant, Theorem 5 demonstrates that any algorithm that does not know $\alpha^{\star}$ will incur a regret that is polynomially larger than the optimal regret attainable by Algorithm 1 with the knowledge of $\alpha^{\star}$. This result on separation shows that batch learning for nonparametric bandits is indeed much harder than the fully online case to some extent, where adaptivity to $\alpha^{\star}$ could be achieved for free [41].
The intuition behind the proof of Theorem 5 is that since the algorithm does not know $\alpha^{\star}$, it has little hope to pick the first batch size $t_{1}$ optimally. If $t_{1}$ is too large, then the adversary can choose a large $\alpha^{\star}$, which corresponds to the family of reward functions with larger gaps, so that the algorithm’s regret during the first batch explodes. If $t_{1}$ is too small, then the adversary can choose a small $\alpha^{\star}$, which corresponds to the family of reward functions with smaller gaps. In this case, the algorithm’s knowledge gathered during the first batch is not enough to distinguish the arms and its regret will explode in later batches.
6 Discussions
@@
Adapting to the margin parameter.
While we have shown that adaptivity to the margin parameter is not possible when the batch constraint is stringent, i.e., when $M$ is constant, it leaves open the question of designing optimal adaptive algorithm when $M$ is large, say $M\asymp\log\log T$. In fact, when $M=T$, i.e., in the fully online setting, [41] provides an adaptively binned successive elimination algorithm that is capable of adapting to the margin parameter optimally.
Gap-dependent regret bounds.
The separation between the arms for nonparametric bandits is controlled by the margin parameter $\alpha$. The current paper focuses on the family of non-degenerate nonparametric bandit instances. When $\alpha\rightarrow\infty$, the two arms will have constant gap over the entire covariate space, meaning the instance is essentially reduced to the multi-armed bandits without covariates. The BaSE policy with geometric grid from [21] can then be applied to obtain a gap-dependent regret bound. Whether one can develop a procedure that bridges these two regimes under the batched setting is an interesting future direction.
@@ Appendix A Proof of Theorem 2
We reuse the family of hard instances $\mathcal{C}_{z_{i}}$ as defined in (4), and define the mixture distribution
@@
where $q_{i}:\Omega_{s_{i}}\rightarrow[0,1]$ is selected so that the above equality holds, and $\mathbb{P}_{\pi,i,f_{\omega}}$ is the distribution of the observations when $f_{\omega}\in\mathcal{C}_{z_{i}}$. It is easy to see $\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)=1$. Let $\{T_{i}\}_{i=1}^{M}$ and $\{z_{i}\}_{i=1}^{M}$ be strictly increasing sequences of positive integers to be specified later. We pause here to state a useful claim regarding the family $\{Q_{i}\}_{i=1}^{M}$ of mixture distributions, namely, they are all close to each other under the total variation distance.
@@
The following lemma ensures that when a sequence of events $A_{i}$’s are properly designed, at least one of them happens with sufficiently large probability under the mixture distribution.
@@
Assume $z_{i}=\lceil(36T_{i-1}(M-k+1)^{2})^{1/(2\beta+d)}\rceil$. Suppose $\{A_{i}\}_{i=k}^{M}$ is a sequence of events that partition the whole probability space where $1\leq k\leq M$. Assume that each $A_{i}$ can be determined by observations up to $T_{i-1}$. Then, there exists $i^{\star}\in\{k,k+1,...,M\}$ such that $Q_{i^{\star}}(A_{i^{\star}})\geq 1/(2(M-k+1))$.
@@
Let $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}$. If $Q_{i}(A_{i})\geq 1/(2(M-k+1))$, then
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}z_{i}^{-\beta(1+\alpha)}(M-k+1)^{-\frac{1+\alpha}{\alpha}}.$
The proofs of the above lemmas are postponed to the end of this section. Now we are ready to establish the desired claim in the theorem. Define $b\asymp T^{(1-\gamma)/(1-\gamma^{M})}$. For each $1\leq i\leq M$, we set $T_{i}=\lfloor b^{(1-\gamma^{i})/(1-\gamma)}\rfloor$, $z_{i}=\lceil(36T_{i-1}M^{2})^{1/(2\beta+d)}\rceil$, and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. Define the event
$A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}.$
Intuitively, $A_{i}$ models the event when the algorithm’s selected grid points $t_{i-1}$ and $t_{i}$ are suboptimal. When $A_{i}$ happens, the goal is to design a problem instance such that using observations up $t_{i-1}$ cannot distinguish the optimal arm and therefore the policy must incur a large regret between $t_{i-1}$ and $t_{i}$. Such $A_{i}$’s satisfy the assumption in Lemma 3 with $k=1$ and we can combine Lemma 3 and Lemma 4 to obtain
@@ A.2 Proof of Lemma 3
For any $k\leq i\leq M$, we have
$|Q_{M}(A_{i})-Q_{i}(A_{i})|\overset{\mathrm{(i)}}{=}|Q_{M}^{T_{i-1}}(A_{i})-Q_{i}^{T_{i-1}}(A_{i})|\overset{\mathrm{(ii)}}{\leq}\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\overset{\mathrm{(iii)}}{\leq}\frac{1}{2(M-k+1)},$ (15)
@@
$\sum_{i=k}^{M}Q_{i}(A_{i})$ $=Q_{M}(A_{M})+\sum_{i=k}^{M-1}Q_{i}(A_{i})$
$=Q_{M}(A_{M})+\sum_{i=k}^{M-1}(Q_{i}(A_{i})-Q_{M}(A_{i})+Q_{M}(A_{i}))$
$\overset{\mathrm{(iv)}}{\geq}Q_{M}(A_{M})+\sum_{i=k}^{M-1}(Q_{M}(A_{i})-\frac{1}{2(M-k+1)})\geq\sum_{i=k}^{M}Q_{M}(A_{i})-\frac{1}{2}\overset{\mathrm{(v)}}{=}\frac{1}{2},$
where step (iv) uses inequality (15), and step (v) uses the fact that $\sum_{i=k}^{M}Q_{M}(A_{i})=1$.
@@ A.3 Proof of Lemma 4
$=\frac{1}{2^{s_{i}}}\cdot\frac{T_{i}}{z_{i}^{d}}\left(s_{i}2^{s_{i}-1}Q_{i}(A_{i})-\frac{3}{2}s_{i}2^{s_{i}-1}\sqrt{\frac{1}{36(M-k+1)^{2}}}\right)$
$\geq\frac{1}{8}T_{i}z_{i}^{-\alpha\beta}\frac{1}{M-k+1},$
where the last inequality uses the assumption $Q_{i}(A_{i})\geq 1/(2(M-k+1))$. Therefore, we arrive at
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}^{-\frac{1}{\alpha}}\left[\sup_{f\in\mathcal{C}_{z_{i}}}S_{t_{i}}(\pi;f)\right]^{\frac{1+\alpha}{\alpha}}\apprge T_{i}z_{i}^{-\beta(1+\alpha)}(M-k+1)^{-\frac{1+\alpha}{\alpha}}.$